A car A1 of mass 30kg moving with a velocity of 2N/s hit another car A2 of mass 50kg moving in the same direction with car A1. If the common velocity of the two cars as they hit each other was 3N/s. Find the velocity of car A2.
1
Expert's answer
2020-05-07T09:49:38-0400
"m_1v_1+m_2v_2=(m_1+m_2)v"

"30(2)+50v_2=(30+50)3"

"v_2=3.6\frac{m}{s}"